40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 527     Entry time: Mon Jun 9 17:57:59 2008
Author: Yoichi 
Type: Configuration 
Category: PSL 
Subject: PMC input power backed to the original 
I rotated back the HWP before the PBS to restore the input laser power to the PMC.
Now the reading of PSL-PMC_PMCTRANSPD is 2.7.
ELOG V3.1.3-